# Coherent errors in stabilizer codes caused by quasistatic phase damping

### v1

@@ I Introduction
Quantum computing [1] promises efficient solutions to technologically relevant computational problems, but the realization of this great potential is hindered by errors in the physical implementation. Quantum error correction offers a potential solution to this by the redundant encoding of a small number of logical qubits in a greater number of physical qubits. The effect of Pauli errors during quantum error correction is well-understood due to the efficient simulability [2] and the existence of mappings to classical disordered spin models [3, 4, 5, 6, 7]. However, the exploration of more realistic error models is considered to be a harder challenge.
@@
In this work, we study the performance of Pauli stabilizer codes in the presence of quasistatic random coherent errors. (i) We identify quantum computing architectures where the effect of $1/f$ noise is well described by an error model we term ‘quasistatic phase damping’, which accounts for quasistatic random coherent errors whose rotation axis (angle) is uniform (non-uniform) across the qubit register. (ii) We show that this quasistatic random coherent error model for Pauli stabilizer codes [11] is equivalent to independent Pauli phase-flip errors in the case of a single cycle of error detection or error correction. For multiple syndrome measurement cycles the model cannot be described with independent phase flips on the physical level. However, it is an effective Pauli model on the logical level. It is known that for single qubit error models stabilizer measurements assymptotically decohers the noise on the logical level [12]. Our results show that for quasistatic phase damping this decoherence process is exact for all code sizes. (iii) We utilize the Fermionic Linear Optics simulation framework based on the Majorana-fermionic representation of the surface code [8], combined with phenomenological readout errors [10], to perform numerical experiments predicting the performance of the surface code in the presence of the quasistatic coherent error model. (iv) We establish a surface code error-correction threshold at a physical error rate $p_{\text{th}}\approx 2.85\%$, including the quasistatic coherent errors and the readout errors. (v) We generate the threshold phase diagram in the plane spanned by the strengths of the coherent error and the readout error. (vi) We discuss the implications of our results for spin qubits and superconducting qubits.
The rest of this paper is structured as follows. In Sec. II, we introduce the error model of quasistatic phase damping, and pinpoint the relevant qubit platforms affected by this type of noise. In Sec. III, we compare and contrast our error model with independent Pauli errors through a minimal example of the 2-qubit repetition code. We generalize this comparison for Pauli stabilizer codes in Sec. IV. In Sec. V and Sec. VI, we present our results on the combined effect of quasistatic phase damping and readout errors on the surface code, and relate these results to semiconductor spin qubits and superconducting qubits.
@@ II Quasistatic phase damping as a model of random coherent errors
In certain cases, the inhomogeneous dephasing time $T_{2}^{*}$ of a superconducting qubit, measured by the Ramsey experiment, is exceeded by the qubit relaxation time $T_{1}$. Furthermore, it is common to find that the Hahn echo improves qubit coherence, i.e., the corresponding decoherence time $T_{2,\text{echo}}$ exceeds $T_{2}^{*}$. The coexistence of these two features, i.e., (i) a dephasing-limited coherence time, and (ii) a functioning Hahn echo, is consistent with, and often interpreted as evidence for, classical electromagnetic noise that induces Larmor-frequency fluctuations. Furthermore, it is often found that the corresponding noise spectrum depending on the frequency $f$ is proportional to $1/f$.
@@
The quasistatic phase damping model is relevant for semiconductor spin qubits as well, since spin qubit decoherence is often dominated by $1/f$ noise. [22, 23] Note that in such devices, $1/f$ noise is attributed either to externally imposed electromagnetic fluctuations, e.g., gate noise, magnet noise, fluctuating charge traps, or to hyperfine interaction, i.e., magnetic noise exerted on the electron spin by the nuclear spins covered by the electron.
@@ III Example: 2-qubit repetition code
In this section, we study quasistatic phase damping using the example of a 2-qubit phase-flip code, as depicted on Fig. 1. We start with the analysis of a single cycle of syndrome measurement in Sec. III.1, and show that in this case, quasistatic phase damping is equivalent to a simple uncorrelated phase-flip channel. In Sec. III.2, explicit calculation of the density matrix reveals that this is not the case for two cycles of error detection. We quantify the difference between our model and the simple Pauli phase-flip model in Sec. III.2.3.
@@
Here, $A$ and $B$ label the two data qubits. The aim of the protocol is to preserve the value of the logical $\hat{Z}$ operators of the code, which reads,
@@
Figure 1: Schematics of the $2$-qubit phase-flip repetition code, data qubits suffering quasistatic phase damping.
@@ III.2.2 Quasistatic phase damping
From this result, we conclude that in this two-cyle case, the effect of quasistatic phase damping cannot be described as random independent phase-flip errors with the same error rate $p$ on both qubits. This is apparent by a comparison of Eqs. (III.2.1) and (III.2.2): for independent homogeneous phase-flip errors, there are only 3 different ${c}_{\boldsymbol{s}}$, ${d}_{\boldsymbol{s}}$ coefficients, while for quasistatic phase damping, there are 4 different coefficients.
@@ IV Quasistatic phase damping in stabilizer codes
In this section, we generalize the results above, i.e., compare the quasistatic phase damping (quasistatic random coherent error) model with the uncorrelated phase-flip model, for Pauli stabilizer codes in general. In Sec. IV.1, we prove that the two models are equivalent in the case of a single cycle of error detection or error correction. In Sec. IV.2, we show that on the logical level, quasistatic phase damping is equivalent to a Pauli channel, however, this is not the one resulting from the simple uncorrelated phase-flip error model in general (see previous section as an example).
@@ IV.1 Single cycle of error detection and error correction
$P(\hat{E})=\prod_{j=1}^{n}\langle\cos^{2}\theta_{j}\rangle_{\boldsymbol{\theta}}^{1-n_{\hat{E}}(j)}\langle\sin^{2}\theta_{j}\rangle_{\boldsymbol{\theta}}^{n_{\hat{E}}(j)}$ (36)
@@
A result analogous to Eq. (35) holds for a single cycle of error correction as well, since the correction operation $\hat{C}_{s}$ depends only on the syndrome $s$. Hence, the state of the code after error correction has the following form:
@@ IV.2 Multiple cycles
We next consider multiple, $t$ cycles of error detection: we apply the independent coherent rotations of the qubits with angles $\theta_{1},\ldots,\theta_{n}$, and the measurements repeatedly. We show that in this case, the average effect of quasistatic phase damping can be represented only with a more complicated Pauli channel.
@@
but they can be negative as well (see Table 1 in the Appendix, depending on which power of the imaginary unit appears in the corresponding terms. Real values are guaranteed by the fact that the terms containing odd powers of $(i\sin\theta_{j})$ will vanish after averaging. In Eq. (41), we only get a non-zero contribution when $\hat{E}(\boldsymbol{E})=\hat{E}(\boldsymbol{E^{\prime}})$.
@@
$\hat{\rho}_{d}(\boldsymbol{s})=\sum_{\boldsymbol{\alpha}}\mathcal{P}_{\boldsymbol{\alpha}}(\boldsymbol{s})\hat{Z}_{\boldsymbol{\alpha}}\hat{E}_{\boldsymbol{s}}\hat{\rho}_{0}\hat{E}_{\boldsymbol{s}}\hat{Z}_{\boldsymbol{\alpha}},$ (46)
@@
$\mathcal{P}_{\boldsymbol{\alpha}}(\boldsymbol{s})=\sum_{\begin{subarray}{c}\boldsymbol{E}\in\mathcal{D}_{\boldsymbol{s}}\\ \hat{E}(\boldsymbol{E})=\hat{Z}_{\boldsymbol{\alpha}}\hat{E}_{\boldsymbol{s}}\end{subarray}}\tilde{P}(\boldsymbol{E}).$ (48)
Furthermore, we defined $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{k})\in\{0,1\}^{k}$, the operator $\hat{Z}^{L}_{j}$ is defined as the logical Z operator acting on the $j$-th logical qubit, and $\hat{E}_{\boldsymbol{s}}$ is an arbitrary merged error string consistent with the syndrome $s$. The $\mathcal{P}_{\boldsymbol{\alpha}}(\boldsymbol{s})$ quantities are syndrome-dependent probabilities (generalizing the coefficients $c_{\boldsymbol{s}}$ and $d_{\boldsymbol{s}}$ from Sec. III.2), which cannot be derived from independent single-qubit Z error probabilities.
Therefore, we conclude that for multiple cycles, the quasistatic phase damping channel is equivalent to an incoherent Pauli channel as written in Eq. (IV.2), as well as an incoherent Pauli channel on the logical level, as written Eq. (46). However, the Pauli channel in Eq. (IV.2) does not describe, in general, independent homogeneous single-qubit phase flips, as we illustrated by the counterexample in Sec. III.2.2.
@@ V Surface code error threshold
In this section, we numerically demonstrate that the surface code in the presence of quasistatic phase damping and readout errors exhibits a threshold behavior. To this end, we apply the Fermionic Linear Optics (FLO) simulation technique [8, 9, 10]. In Sec. V.1 we briefly summarize the basics of the surface code and the effect of coherent errors. In Sec. V.2 we discuss error correction in the presence of readout errors. Finally, we present our numerical results in Sec. V.3.
@@ V.3 Numerical results
Fig. 4 shows that by increasing the code distance, we observe threshold behavior. When the rates of physical and readout errors are the same, $p=q=\langle\sin^{2}\theta\rangle$, we find that the threshold is $p_{\text{th}}\approx 2.85\%$. This is close to the threshold with Pauli Z errors and readout errors, which for the toric code is $2.93\pm 0.02\%$ [4]. We also carried out Monte Carlo simulations to investigate phase-flip errors. Interestingly, we have found that the threshold value coincides with the above mentioned $2.85\%$, however, the corresponding logical error rate is different for the two error models: in our case $p_{L}(p_{\text{th}})\approx 7\%$, but for the combination of phase-flip and readout errors it is slightly higher, around $8.5\%$. This relative difference is sustained throughout the physical error rate window of Fig. 4 (not shown).
Figure 4: Surface code threshold with quasistatic phase damping and readout errors for $p=q$. Data points show the numerically obtained logical error rate (maximum infidelity), for different code distances $d$, with $p=q=\langle\sin^{2}\theta\rangle$ being the physical error rate $p$ and readout error rate $q$, that are kept equal for this plot. Each point is a result of $N_{\text{sample}}=10,000\times d\times 100$ simulations: we sampled the multi-cycle syndrome $\boldsymbol{s}$ for $10,000$ random $\boldsymbol{\theta}$ angles, simulating $100$ random readout error configurations in each case. The graph confirms the existence of a threshold at $p_{\text{th}}\approx 2.85\%$.
We have also investigated how varying the rate $p$ of random coherent errors and $q$ of readout errors independently affect the threshold of the surface code. Results are shown in 5. For many pairs of $p$ and $q$ we numerically assessed whether scaling up the surface code decreases the logical error rate (scalable QEC, green area) or increases it (unscalable QEC, red area). The threshold should be in between these regions. The results shown in Fig. 5 show that the surface code is more sensitive to quasistatic phase damping than to readout errors. If the physical error rate is at the $1\%$ level, the surface code is quite robust against readout errors, scalable even with relatively high $q\approx 8\%$. However, if the readout error rate is at the $1\%$ level, the surface code still requires the coherent error rate to be below $4\%$.

### v3

@@ I Introduction
Quantum computing [1] promises efficient solutions to technologically relevant computational problems, but the realization of this great potential is hindered by errors in the physical implementation. Quantum error correction, with a thirty-year history from the fundamentals [2, 1] to the most recent experimental milestones [3, 4, 5, 6, 7, 8, 9], offers a potential solution to this by the redundant encoding of a small number of logical qubits in a greater number of physical qubits. The effect of Pauli errors during quantum error correction is well-understood due to the efficient simulability [10] and the existence of mappings to classical disordered spin models [11, 12, 13, 14, 15]. However, the exploration of more realistic error models is an ever-present challenge.
@@
In this work, we study the performance of Pauli stabilizer codes against quasistatic phase damping. In this error model, physical qubits suffer coherent Z-rotations, with rotation angle randomly varying from qubit to qubit, but timewise constant throughout the entire quantum circuit; however, for each repetition of the circuit (shot) a new random value for each angle is chosen. (i) We identify quantum computing architectures where this error model is relevant, and describes $1/f$ noise. (ii) For Pauli stabilizer codes [19] we analyze how this error model is related to independent Pauli phase-flip errors. For a single cycle of error detection or error correction, we find the two error models are equivalent. For multiple syndrome measurement cycles this is no longer true, however, quasistatic phase damping is an effective – correlated – Pauli model on the logical level. It is known that for single qubit error models stabilizer measurements asymptotically decohere the noise on the logical level [20]. Our results show that for quasistatic phase damping this decoherence process is exact for all code sizes. (iii) We utilize the Fermionic Linear Optics simulation framework based on the Majorana-fermionic representation of the surface code [16], combined with phenomenological readout errors [18], to perform numerical experiments predicting the performance of the surface code in the presence of the quasistatic coherent error model. (iv) We establish a surface code error-correction threshold at a physical error rate $p_{\text{th}}\approx 2.85\%$, including the quasistatic coherent errors and the readout errors. (v) We generate the threshold phase diagram in the plane spanned by the strengths of the coherent error and the readout error. (vi) We discuss the implications of our results for spin qubits and superconducting qubits.
The rest of this paper is structured as follows. In Sec. II, we introduce the error model of quasistatic phase damping and pinpoint the relevant qubit platforms affected by this type of noise. In Sec. III, we compare and contrast our error model with independent Pauli errors through a minimal example of the 2-qubit phase-flip repetition code. We generalize this comparison for Pauli stabilizer codes in Sec. IV. In Sec. V and Sec. VI, we present our results on the combined effect of quasistatic phase damping and readout errors on the surface code, and relate these results to semiconductor spin qubits and superconducting qubits.
@@ II Quasistatic phase damping as a model of random coherent errors
In certain cases, the inhomogeneous dephasing time $T_{2}^{*}$ of a superconducting qubit, measured by the Ramsey experiment, is exceeded by the qubit relaxation time $T_{1}$; furthermore, it is common to find that the Hahn echo improves qubit coherence, i.e., the corresponding decoherence time $T_{2,\text{echo}}$ exceeds $T_{2}^{*}$ [29, 30, 31]. The coexistence of these two features, i.e., (i) a dephasing-limited coherence time, and (ii) a functioning Hahn echo, is consistent with, and often interpreted as evidence for, classical electromagnetic noise that induces Larmor-frequency fluctuations. Furthermore, it is often found that the corresponding noise spectrum depending on the frequency $f$ is proportional to $1/f$.
@@
The quasistatic phase damping model is relevant for semiconductor spin qubits as well, since spin qubit decoherence is often dominated by $1/f$ noise. [33, 34, 35, 36] Note that in such devices, $1/f$ noise is attributed either to externally imposed electromagnetic fluctuations, e.g., gate noise, magnet noise, fluctuating charge traps, or to hyperfine interaction, i.e., magnetic noise exerted on the electron spin by the nuclear spins covered by the electron.
@@ III Example: 2-qubit repetition code
In this section, we study quasistatic phase damping using the example of a 2-qubit phase-flip code, as depicted on Fig. 1. We start with the analysis of a single cycle of syndrome measurement in Sec. III.1, and show that in this case, quasistatic phase damping is equivalent to a simple uncorrelated phase-flip channel. In Sec. III.2, explicit calculation of the density matrix reveals that this is not the case for two cycles of error detection. We quantify the difference between our model and the simple Pauli phase-flip model in Sec. III.2.3. This simple example proves that quasistatic phase damping is a type of error that is distinct from phase-flip Pauli errors. This observation motivates the study of the stabiliser codes subject to quasistatic phase damping, which we do, focusing on the surface code, in Secs. V and VI.
@@
Here, $A$ and $B$ label the two data qubits. The protocol aims to preserve the value of the logical $\hat{Z}$ operators of the code, which reads,
@@
Figure 1: Schematic circuit representation of the $2$-qubit phase-flip repetition code, data qubits A and B suffering quasistatic phase damping. The figure shows elements of two subsequent cycles in a multi-cycle error detection procedure. The gates labelled as $R_{z}(\theta_{A,B})$ represent coherent errors.
@@ III.2.2 Quasistatic phase damping
From this result, we conclude that in this two-cyle case, the effect of quasistatic phase damping cannot be described as random independent phase-flip errors with the same error rate $p$ on both qubits. This is apparent by a comparison of Eqs. (21) and (22): for independent homogeneous phase-flip errors, there are only 3 different ${c}_{\boldsymbol{s}}$, ${d}_{\boldsymbol{s}}$ coefficients, while for quasistatic phase damping, there are 4 different coefficients.
@@ IV Quasistatic phase damping in stabilizer codes
In this section, we generalize the results above, i.e., compare the quasistatic phase damping (quasistatic random coherent error) model with the uncorrelated phase-flip model, for Pauli stabilizer codes in general. In Sec. IV.1, we prove that the two models are equivalent in the case of a single cycle of error detection or error correction. In Sec. IV.2, we show that on the logical level, quasistatic phase damping is equivalent to a Pauli channel, however, this is not the one resulting from the simple uncorrelated phase-flip error model in general (see previous section as an example). Motivated by this difference of the error models, we will study the effect of quasistatic phase damping on the surface code in Secs. V and VI.
@@ IV.1 Single cycle of error detection and error correction
$P(\hat{E})=\prod_{j=1}^{n}\langle\cos^{2}\theta_{j}\rangle_{\boldsymbol{\theta}}^{1-n_{\hat{E}}(j)}\langle\sin^{2}\theta_{j}\rangle_{\boldsymbol{\theta}}^{n_{\hat{E}}(j)}.$ (36)
@@
A result analogous to Eq. (35) holds for a single cycle of error correction as well since the correction operation $\hat{C}_{s}$ depends only on the syndrome $s$. Hence, the state of the code after error correction has the following form:
@@ IV.2 Multiple cycles
We next consider multiple, $t$ cycles of error detection: we apply the independent coherent rotations of the qubits with angles $\theta_{1},\ldots,\theta_{n}$, and the measurements repeatedly. We show that in this case, the average effect of quasistatic phase damping can be represented only with a more complicated Pauli channel. It is important to note that the angles $\theta_{j}$ do not change from cycle to cycle: if they did, this error model would – after averaging over the random angles – be equivalent to independent phase-flip errors.
@@
but they can be negative as well (see Table 1 in the Appendix), depending on which power of the imaginary unit appears in the corresponding terms. Real values are guaranteed by the fact that the terms containing odd powers of $(i\sin\theta_{j})$ will vanish after averaging. In Eq. (41), we only get a non-zero contribution when $\hat{E}(\boldsymbol{E})=\hat{E}(\boldsymbol{E^{\prime}})$.
@@
$\hat{\rho}_{t}(\boldsymbol{s})=\sum_{\boldsymbol{\alpha}}\tilde{\mathcal{P}}_{\boldsymbol{\alpha}}(\boldsymbol{s})\hat{Z}_{\boldsymbol{\alpha}}\hat{E}_{\boldsymbol{s}}\hat{\rho}_{0}\hat{E}_{\boldsymbol{s}}\hat{Z}_{\boldsymbol{\alpha}},$ (44)
@@
$\tilde{\mathcal{P}}_{\boldsymbol{\alpha}}(\boldsymbol{s})=\sum_{\begin{subarray}{c}\boldsymbol{E}\in\mathcal{D}_{\boldsymbol{s}}\\ \hat{E}(\boldsymbol{E})=\hat{Z}_{\boldsymbol{\alpha}}\hat{E}_{\boldsymbol{s}}\end{subarray}}\tilde{P}(\boldsymbol{E}).$ (46)
Furthermore, we defined $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{k})\in\{0,1\}^{k}$, the operator $\hat{Z}^{L}_{j}$ is defined as the logical Z operator acting on the $j$-th logical qubit, and $\hat{E}_{\boldsymbol{s}}$ is an arbitrary merged error string consistent with the syndrome $s$. The $\tilde{\mathcal{P}}_{\boldsymbol{\alpha}}(\boldsymbol{s})$ quantities are syndrome-dependent probabilities, meaning that they are positive (cf. Eq. (8.152) of [1]), and satisfy
$\sum_{\boldsymbol{\alpha},\boldsymbol{s}}\tilde{\mathcal{P}}_{\boldsymbol{\alpha}}(\boldsymbol{s})=1.$ (47)
These probabilities are the generalization of the single-logical-qubit coefficients $\tilde{c}_{\boldsymbol{s}}$ and $\tilde{d}_{\boldsymbol{s}}$ of Eq. (22), e.g. $\tilde{c}_{++}=\tilde{\mathcal{P}}_{0}(++)$, $\tilde{d}_{++}=\tilde{\mathcal{P}}_{1}(++)$. As discussed in Sec. III.2, these probabilities cannot be derived from independent single-qubit Z error probabilities.
We have seen that Eq. (IV.2) is not a well-defined Pauli channel on the physical level, because the $\tilde{P}(\boldsymbol{E})$ weights can be negative, thus these are not probabilities. However, we can rephrase Eq. (IV.2) by a new set of amplitudes $\tilde{P}^{\prime}(\boldsymbol{E})$ that are positive, and hence define a proper Pauli channel on the physical level as well. This new set of amplitudes can be defined by redistributing the amplitudes within each class of cycle-resolved error scenarios ($\boldsymbol{E}$), where the classes are indexed by $\boldsymbol{s}$ and $\boldsymbol{\alpha}$, and are defined via $\boldsymbol{E}\in\mathcal{D}_{\boldsymbol{s}}$, $\hat{E}(\boldsymbol{E})=\hat{Z}_{\boldsymbol{\alpha}}\hat{E}_{\boldsymbol{s}}$. In words, two cycle-resolved error scenarios belong to the same class, if they imply the same syndrome and the same merged error string. By performing such a redistribution $\tilde{P}\to\tilde{P}^{\prime}$ of amplitudes, the error channel in Eq. (44) does not change: the $\tilde{\mathcal{P}}_{\boldsymbol{\alpha}}(\boldsymbol{s})$ probabilities remain the same. Note that this redistribution is not unique; one way to do it is to choose one representative cycle-resolved error scenario $\boldsymbol{E}^{\prime}_{\boldsymbol{s},\boldsymbol{\alpha}}$ from each error class, and define new amplitudes, which are in fact non-negative, as:
$\tilde{P}^{\prime}(\boldsymbol{E}^{\prime}_{\boldsymbol{s},\boldsymbol{\alpha}})=\tilde{\mathcal{P}}_{\boldsymbol{\alpha}}(\boldsymbol{s}),$ (48)
and $\tilde{P}^{\prime}(\boldsymbol{E})=0$ for all $\boldsymbol{E}$ that is not in the set of $\boldsymbol{E}^{\prime}_{\boldsymbol{s},\boldsymbol{\alpha}}$ representatives. This shows that our error model can be described as a Pauli model on the physical level. However, this Pauli model is highly correlated. It is an open question whether it is possible to construct an error model with local (in both space and time) correlations by redefining these weights differently. This question is beyond the scope of this work.
We conclude that for multiple cycles, the quasistatic phase damping channel is equivalent to an incoherent Pauli channel as written in Eq. (IV.2), as well as an incoherent Pauli channel on the logical level, as written Eq. (44). However, the Pauli channel in Eq. (IV.2) does not describe, in general, independent homogeneous single-qubit phase flips, as we illustrated by the counterexample in Sec. III.2.2.
@@ V Surface code error threshold
So far, we have shown that for stabiliser codes applied for multi-cycle error correction, the effect of quasistatic phase damping is distinct from that of phase-flip Pauli errors. This motivates studies of the noise resilience in this new setting. In this section, we do this by focusing on the surface code. In particular, we numerically demonstrate that the surface code in the presence of quasistatic phase damping and readout errors exhibits a threshold behavior. To this end, we apply the Fermionic Linear Optics (FLO) simulation technique [16, 17, 18]. In Sec. V.1 we briefly summarize the basics of the surface code and the effect of coherent errors. In Sec. V.2 we discuss error correction in the presence of readout errors. Finally, we present our numerical results in Sec. V.3, showing that the threshold for quasistatic phase damping is, up to our numerical precision, the same as the threshold for Pauli phase flips, whereas the logical error rate at threshold is smaller in the the former case (7%) than in the latter case (8.5%).
@@ V.3 Numerical results
Fig. 4a shows that by increasing the code distance, we observe threshold behavior. When the rates of physical and readout errors are the same, $p=q=\langle\sin^{2}\theta\rangle$, we find that the threshold is $p_{\text{th}}\approx 2.85\%$ (see Fig. 4b). This is close to the threshold with Pauli Z errors and readout errors, which for the toric code is $2.93\pm 0.02\%$ [12]. We also carried out Monte Carlo simulations to investigate phase-flip errors. Interestingly, we have found that the threshold value coincides with the above mentioned $2.85\%$, however, the corresponding logical error rate is different for the two error models: in our case $p_{L}(p_{\text{th}})\approx 7\%$, but for the combination of phase-flip and readout errors it is slightly higher, around $8.5\%$. This relative difference is sustained throughout the physical error rate window of Fig. 4b (not shown).
Figure 4: Surface code threshold with quasistatic phase damping and readout errors for $p=q$. Data points show the numerically obtained logical error rate (maximum infidelity), for different code distances $d$, with $p=q=\langle\sin^{2}\theta\rangle$ being the physical error rate $p$ and readout error rate $q$, that are kept equal for this plot. Each point is a result of $N_{\text{sample}}=10,000\times d\times 100$ simulations: we sampled the multi-cycle syndrome $\boldsymbol{s}$ for $10,000$ random $\boldsymbol{\theta}$ angles, simulating $100$ random readout error configurations in each case. (a) Logical error rate for a wide range of physical error rates. (b) Zoom-in of (a) into a narrow window of physical error rates in the vicinity of the threshold. The graph confirms the existence of a threshold at $p_{\text{th}}\approx 2.85\%$.
We have also investigated how varying the rate $p$ of random coherent errors and $q$ of readout errors independently affect the threshold of the surface code. Results are shown in 5. For many pairs of $p$ and $q$ we numerically assessed whether scaling up the surface code decreases the logical error rate (scalable quantum error correction, green area) or increases it (unscalable quantum error correction, red area). The threshold should be in between these regions. For example, if the physical error rate is at the $1\%$ level, the surface code is quite robust against readout errors, scalable even with relatively high $q\approx 8\%$. However, if the readout error rate is at the $1\%$ level, the surface code requires the coherent error rate to be below $4\%$.